### v1

# Towards Generating Executable Metamorphic Relations Using Large Language Models

Abstract
Metamorphic testing (MT) has proven to be a successful solution to automating testing and addressing the oracle problem. However, it entails manually deriving metamorphic relations (MRs) and converting them into an executable form; these steps are time-consuming and may prevent the adoption of MT. In this paper, we propose an approach for automatically deriving executable MRs (EMRs) from requirements using large language models (LLMs). Instead of merely asking the LLM to produce EMRs, our approach relies on a few-shot prompting strategy to instruct the LLM to perform activities in the MT process, by providing requirements and API specifications, as one would do with software engineers. To assess the feasibility of our approach, we conducted a questionnaire-based survey in collaboration with Siemens Industry Software, focusing on four of their software applications. Additionally, we evaluated the accuracy of the generated EMRs for a web application. The outcomes of our study are highly promising, as they demonstrate the capability of our approach to generate MRs and EMRs that are both comprehensible and pertinent for testing purposes.
Index Terms:
metamorphic testing, large language model, executable metamorphic relations
I Introduction
In many sectors, software is typically verified with executable test cases that, at a high level, consist of a set of inputs provided to the software under test (SUT) and a set of test assertions verifying that the SUT outputs match the expected results. However, defining executable test cases is expensive because of the many intellectual activities involved. This cost limits the number of test cases that can be implemented. Moreover, software faults are often subtle and triggered by a narrow portion of the input domain. Therefore, they are detected only after exercising the SUT with a large set of test inputs.
Although several tools for the automated generation of test inputs had been developed [1, 2], they can only identify crashes or regressions faults, because they lack the capability of determining what the expected output for the software should be. Further, manually specifying test assertions for thousands of automatically generated test inputs is practically infeasible. The impossibility to programmatically derive expected outputs is known as the oracle problem [3].
Metamorphic testing (MT) has recently gained success as a solution to address the oracle problem in many contexts [4]. MT relies on the concept of metamorphic relations (MRs), which describe the relationships between input transformations and expected output changes, serving as oracles in software testing. Briefly, to perform MT on the SUT, engineers first need to derive MRs from the SUT’s requirements and then convert the derived MRs into an executable form to determine the test outcome. These steps require substantial manual effort, which increases the MT cost and may prevent its adoption.
State of the art. Many MT approaches have been developed for testing various SUTs, such as search engines [5], web applications [6], and embedded systems [7], accounting for the requirements specific to these SUTs. However, these approaches heavily rely on manual efforts to elicit MRs, inherently limiting the applicability of MT. To reduce the cost of defining MRs, researchers have proposed relying on meta-heuristic search [8, 9] and genetic programming [10] to automatically derive MRs from execution data. Although promising, such approaches require several executions of the SUT, which makes them suitable for small programs, but may not scale for large software where executions may take several minutes and a lot of outputs are produced.
Recently, a few studies [11, 12] have proposed the use of LLMs to automatically derive MRs from the knowledge base of LLMs. However, these studies are limited to deriving MRs for SUTs already known in the adopted LLMs. In other words, there are no studies demonstrating the applicability of LLMs to derive MRs that account for the requirements specific to an SUT that was unseen during the LLM training phases. Considering that, in industry, software testing activities often target new products implementing requirements different from those implemented by existing systems, LLM-based approaches for the automated generation of MRs should be capable of handling unseen requirements. This capability is what we aim to demonstrate in this paper. Additionally, we note that, to the best of our knowledge, there is no work that attempts to automate the conversion of MRs into an executable form, which is required to fully automate the MT process. Chaleshtari et al. [6] proposed the adoption of a domain-specific language (DSL) to automate the execution of MRs; however, DSL-based MRs may have limited readability, contrary to MRs in natural language. In summary, there is no solution in the literature that aims at fully automating the MT process, involving the derivation of MRs from SUT-specific requirements and the generation of executable MRs (EMRs), for the automated determination of test results in MT. An alternative is the automated generation of test cases—including oracles—using LLMs, which is under active development [13]. However, since LLM-generated code may still require human validation, we believe that validating EMRs is more cost-effective than validating test cases. Indeed, a single validated EMR enables exercising the SUT with multiple inputs, which might otherwise be exercised by different test cases, each needing validation.
Contribution. In this paper, we propose an approach for automatically deriving EMRs from the SUT’s requirements using large language models (LLMs). Differently from existing work, our approach does not merely involve querying an LLM for MRs based solely on the LLM knowledge base. Instead, we rely on prompt engineering practices to teach the LLM both (1) the specifications of the SUT, which are necessary to derive MRs, and (2) the DSL to use for EMRs, which is necessary to enable MR execution.
We evaluated the feasibility of our approach through two experiments: (1) by conducting a questionnaire-based survey in collaboration with Siemens Industry Software (SISW) involving four of their software applications, and (2) by assessing the correctness of the generated EMRs for a web application. In our experiments, we used the web interface of ChatGPT [14] to access the underlying LLM, and provide prompts and generate responses for deriving MRs and subsequently converting them into EMRs. When converting MRs into EMRs, we instructed ChatGPT to use SMRL [15] as a DSL for specifying MRs in an executable form. The results are promising, indicating that the generated MRs and EMRs are understandable and relevant for testing, and that the EMRs were correctly converted from their corresponding MRs.
II Approach
II-A Overview
Figure 1: Our approach for generating EMRs using LLMs.
Our approach, shown in Figure 1, advances the metamorphic testing (MT) process by introducing two automated activities: “derive MRs” and “generate EMRs”. The “derive MRs” activity takes as input requirements specifications and outputs MRs in natural language. Unlike existing work that queries an LLM for MRs based on the LLM built-in knowledge, we provide SUT specification documents to the LLM. This enables our approach to derive SUT-specific MRs, which would otherwise be unattainable because the LLM lacks awareness of the specific software properties that need to be considered. The “generate EMRs” activity takes as input MRs and the SUT (depicted as “software system” in Fig. 1). It outputs EMRs, which are executable programs used to verify that the SUT satisfies the properties captured by the MR. In the subsequent subsections, we further describe these activities, focusing on their implementation as adopted in our preliminary experiments (see section III).
II-B Deriving metamorphic relations
We aim at automatically deriving MRs from the provided requirements specifications. To this end, our approach relies on ChatGPT and develops a prompt sequence for automation. Since deriving MRs from requirements is a complex task, when using ChatGPT for this purpose, we break the task down into simpler, decomposed steps [^1] We note that when we passed a single, monolithic query to ChatGPT for deriving MRs, it was unable to derive meaningful MRs .
Our prompt sequence begins by setting the context, i.e., deriving MRs, to guide ChatGPT in the intended direction for subsequent interactions. The remaining sequence, which decomposes the task of deriving MRs, consists of prompts directing ChatGPT to perform simpler steps: (1) read a requirements document, (2) identify sentences relevant to system inputs and outputs, and (3) rewrite these sentences into MRs.
II-C Generating executable metamorphic relations
To automatically generate EMRs, we convert MRs written in natural language using SMRL [15], a DSL developed for specifying EMRs. SMRL provides MR-specific language constructs, which are built on Java, to specify EMRs. Specifically, these constructs enable to specify EMRs in terms of inputs, input transformations, outputs, output changes, and logical expressions to define the expected relations among them. In addition, SMRL is already integrated into the MST-wi framework [6], which generates executable Java code from specified EMRs to automatically perform MT.
Our approach involves instructing ChatGPT to understand SMRL and to use the latter for converting MRs into EMRs. Specifically, when using ChatGPT, we first set the context by indicating that the subsequent interactions will aim at converting MRs into EMRs. We then apply the few-shot learning strategy [16] to instruct ChatGPT to understand SMRL, by providing three pairs of MR and corresponding EMR from prior work [6]. The few-shot learning strategy is suitable in our study context, as SMRL is a state-of-the-art DSL with a limited number of examples available, and it is built on Java, which ChatGPT already knows. In addition, we further instruct ChatGPT to understand the SUT’s APIs by providing the documentation and signature of the API methods (38) used in the EMRs manually implemented in prior work. We then ask ChatGPT to transform MRs into EMRs using SMRL and to invoke the APIs in the EMRs for testing the SUT.
III Experiment Design
To assess the feasibility of our approach, we have carried out two experiments, namely EXP1 and EXP2.
EXP1 - SISW applications
To assess whether our approach practically helps software practitioners, it is important to collect and analyze their perceptions on our work. In collaboration with SISW, we have conducted a questionnaire-based survey to gather their feedback on the MRs and EMRs obtained by applying our approach to their software applications. Due to confidentiality reasons, SISW was not able to share the requirements and the APIs of their applications; instead, SISW selected, based on their interests, four non-confidential technical documents that describe their applications for marine design, wind turbine, aircraft propulsion, and aero-acoustics. They provided us with these documents (83 pages, in total) as requirements specifications. In absence of appropriate APIs, ChatGPT suggests invoking methods that should be implemented to automate the execution of the MR. This emulates the adoption of the approach in early development stages, as in test-driven development, and, therefore, justifies the four selected case study subjects. Further, despite the lack of APIs, considering the preliminary nature of this work, these subjects still enable us to gather feedback from practitioners on how well ChatGPT can generate EMRs. In EXP1, we relied on GPT-4 because it can process documents.
The questionnaire for MRs (resp. EMRs) contains three statements for each MR (resp. EMR). It requests practitioners to indicate their level of agreement with the statements on a Likert scale (i.e., strongly agree, agree, neutral, disagree, and strongly disagree). For example, the questionnaire for MRs contains the following statements: (S1) “The MR is easy to understand”, (S2) “The MR captures a property that needs to be considered when testing the system”, and (S3) “The MR helps identify the expected outputs for given inputs”. These statements, including those for EMRs available online [17], are designed to collect practitioners’ feedback on various aspects, such as complexity, compatibility, and applicability.
Our approach derived 50 MRs and their corresponding EMRs from the four documents. Of these MRs, SISW independently assigned two practitioners to evaluate the 14 MRs derived for the aero-acoustics application, while one was assigned to 36 MRs for the remaining three applications, totaling 64 MRs analyzed. For the EMRs, one practitioner was assigned to all the applications.
EXP2 - Web application
EXP2 evaluates the correctness of the EMRs generated by our approach. Recall that EXP1 studies EMRs that are incomplete with regard to invoking the SUTs’ APIs. Therefore, EXP2 employs a different application, for which we have its API specifications. EXP2 uses the experiment dataset provided by the prior work on MST-wi [6], containing both the APIs enabling the testing of a web application (hereafter, SUT) and MRs that are designed to detect security vulnerabilities. Specifically, in EXP2, we randomly selected ten MRs among the ones published online after the cutoff date for ChatGPT (in EXP2, we relied on GPT-3.5 for this reason). Given the ten MRs and the SUT’s APIs, we applied our approach to produce ten corresponding EMRs.
To assess the correctness of the ten generated EMRs, we inspected and annotated each line of the EMRs with one of the 13 labels we defined, which are available online [17]. These labels are classified into two groups: correct and incorrect. The former contains three labels indicating whether a statement in an EMR uses SMRL constructs correctly, implements part of the corresponding MR exactly, or implements valid alternatives. The latter includes the remaining ten labels, each indicating a different type of error, such as “incorrect use of SMRL constructs” and “misuse of the SUT’s APIs”.
IV Experiment results
EXP1 Results. Table I shows the results of EXP1, focusing on the combined number of responses (i.e., Likert ratings) for each statement across the 64 MRs reviewed by the three practitioners at SISW. From the results of statement 1, we found that in 49 out of the 64 MRs (representing 77%), the practitioners were able to understand the presented MRs and expressed positive feedback, i.e., “strongly agree” or “agree”. In 12 MRs, the practitioners were neutral, and in the remaining 3 MRs, they disagreed with statement 1. Regarding statement 2, the practitioners agreed in 41 out of the 64 MRs (64%), that the MRs capture properties that need to be considered when testing the SUT. In 10 MRs, the practitioners were neutral; and in the remaining 13 MRs, the practitioners disagreed or strongly disagreed with statement 2. In contrast to these results, which are more positive, the responses to statement 3 showed mixed ratings. This statement indicates whether the presented MR helps identify the expected outputs for given inputs. In 18 MRs, the practitioners agreed; in 19 MRs, they were neutral; and in the remaining 27, they disagreed or strongly disagreed.
TABLE I: Responses to the MRs survey, for each statement (S1, S2, S3).
strongly agree
agree
neutral
disagree
strongly disagree
S1
3
46
12
3
0
S2
0
41
10
12
1
S3
0
18
19
24
3
Regarding the non-positive responses to statement 3, we identified possible reasons from the practitioners’ qualitative feedback. In particular, one practitioner, who provided 11 ratings as “disagree” and one as “strongly disagree” to the statement, stated that while most of the MRs are correct and useful for reminding someone what to test, they are too generic to aid in deriving appropriate testing. This aligns with feedback from another practitioner at SISW who, despite not assessing the MRs, collected and reviewed the responses to the MRs questionnaire. The practitioner provided a possible reason for some of the MRs being very generic, stating that it could be attributed to the nature of the documents provided by SISW.
Regarding the survey results for the 50 EMRs in EXP1, one practitioner provided consistent responses across all EMRs, as follows: First, the practitioner perceived that the EMRs are clean and understandable. Second, the practitioner found that the EMRs are consistent with their corresponding MRs. However, the practitioner deemed that not all functions used in the EMRs are feasible to be implemented. Regarding the last opinion, the practitioner noted that the functions in the EMRs, generated by ChatGPT, are specific to each EMR. Therefore, there is little opportunity for reusing these functions across different EMRs, meaning the effort required for implementation is significant. In addition, the specific constraints of the testing environment for their applications may pose challenges in using the exposed APIs to implement functions in the EMRs. Even though EXP1 generated incomplete EMRs with regard to using the SUT’s APIs, as discussed above, this finding suggests that the characteristics of the SUT, such as its APIs, should be considered when further elaborating our approach to deriving EMRs.
EXP2 Results. In EXP2, we annotated each line of the ten EMRs, which were converted from the corresponding ten MRs, for the web application studied in the prior work of MST-wi. These EMRs have a minimum of 10 statements, a mean of 13.8 statements, and a maximum of 20 statements, totaling 138 statements.
Out of the 138 statements in the EMR code, the majority (107 statements) were classified as correctly converted. Specifically, 52 of the 107 statements correctly and exactly implement part of the corresponding MRs. 54 of the 107 statements correctly use SMRL constructs. The remaining 1 statement is valid and correct with respect to the corresponding MR. As for the 31 statements that were deemed incorrect, they can be further categorized as follows: (1) 13 statements incorrectly use APIs; (2) 10 statements incorrectly use SMRL constructs;(3) 6 statements incorrectly implement what is suggested in the corresponding explanations; (4) 2 statements miss implementing what described in the corresponding explanations. Note that each EMR statement is associated with an explanation in natural language produced by ChatGPT.
V Conclusion and future work
In this paper, we have introduced our approach for automatically deriving EMRs from requirements using LLMs. In our preliminary experiments, we applied our approach to four SISW applications and conducted a questionnaire-based survey study to collect opinions from SISW practitioners. In addition, we evaluated the correctness of the EMRs generated by our approach, by applying it to a web application.
In our future work, we plan to explore several key areas related to our current research. A primary focus will be on automating prompt engineering to facilitate the generation of prompts for deriving MRs from requirements and their subsequent conversion into EMRs. Recognizing the significance of input quality, particularly in requirements, another important area will involve developing methods to assist engineers in providing high-quality inputs.
Data availability. Our experiment package with prompts, MRs, EMRs, and questionnaires is available online [17].